A girl throws a ball vertically upward at 10m/s from a roof of a building 20m high.


(a) How long will it take the ball to reach the ground?


(b) What will it's velocity be just before it strikes the ground?

1
Expert's answer
2022-06-06T08:53:49-0400
"y=y_0+v_0t-gt^2\/2"

"y=20+10t-4.9t^2"

(a) at the ground "y=0"

"20+10t-4.9t^2=0"

"t=3.3\:\rm s"

(b)

"v=\sqrt{v_0^2+2gh}=\sqrt{10^2+2*9.8*20}=22\:\rm m/s"